Two waves travelling in the same direction are given by: y1(x,t) = 0.2 sin⁡(3πx-20t+π/2) and y2(x,t) = 0.2 sin⁡(3πx-20t), where x and y are in meters and t is in seconds. The wave function of the resultant wave is:

1
Expert's answer
2021-12-13T16:55:23-0500

"y_1=0.2\sin(3\pi x-20t+\frac{\pi}2),"

"y_2=0.2\sin(3\pi x-20t),"

"y=y_1+y_2=2\cdot 0.2\cos\frac{\pi}4\sin(3\pi x-20t+\frac{\pi}4)=0.28\sin(3\pi x-20t+\frac{\pi}4)."